OpenStudy (anonymous):

Carlotta spent $35 at the market. This was seven dollars less than three times what she spent at the bookstore; how much did she spend there?

OpenStudy (anonymous):

Will fan and medal

OpenStudy (vishweshshrimali5):

Okay so for time being let's assume that it was not Carlotta that spent that money but you... is that fine ? :D

OpenStudy (anonymous):

yep :)

OpenStudy (vishweshshrimali5):

Now, you are on a trip to some place.. I have never visited America so I will go with Washington DC... Someone asked you to visit the market... I know that you have spent $ 35 at the market..

OpenStudy (vishweshshrimali5):

You also spent some money at bookstore... but I don't know how much.. so let's assume that you spent say $ x ... is that fine?

OpenStudy (anonymous):

Yep!

OpenStudy (vishweshshrimali5):

Great! Now, someone told me that "$ 35 was seven dollars less than three times what she spent at the bookstore"... So, let's replace words with numbers...

OpenStudy (vishweshshrimali5):

Okay so 35 dollars will become the number 35... hmm.. what's next.."was seven dollars less that three times what she spent at the bookstore" ... what she spent at bookstore = x... three times of that ... = 3 * x seven dollars less = 3 * x - 7 Now, this whole was 35 dollars.. so So, 3 * x - 7 = 35

OpenStudy (vishweshshrimali5):

Any doubts?

OpenStudy (anonymous):

nope

OpenStudy (vishweshshrimali5):

Great! So, now you only need to solve this equation.. 3x - 7 = 35...can you do this?

OpenStudy (anonymous):

Yeah! Thank you so much :D

OpenStudy (vishweshshrimali5):

:D Great! Your welcome :)